1) An electrolysis experiment is performed according to the equation: 2H20 → 2H₂ + O2

If 4g of water are consumed and 1.3 g of oxygen are recovered, how many grams of hydrogen were produced? What law is being expressed?

2) Silicon dioxide can be purified using carbon to produce silicon and carbon monoxide according to the equation: SiO₂ + 2C → 2CO + Si

If 29 kg of silicon are harvested along with 1.7 kg of CO and 15 kg of carbon were used, how much SiO₂ was used as well?
1
Expert's answer
2018-07-11T07:31:32-0400
1) m(H2O) = m(H2) + m(O2)

m(H2) = 4 g – 1.3 g = 2.7 g

The law of conservation of mass is being expressed.

2) m(SiO2) + m(C) = m(CO) + m(Si)

m(SiO2) + 15 kg = 1.7 kg + 29 kg

m(SiO2) = 30.7 kg – 15 kg = 15.7 kg